OpenStudy (anonymous):

f(x) = √(2x) , then f'(2) = _____ ?

OpenStudy (hba):

@Loushe Do you know how to differentiate ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I got (√2)/2

OpenStudy (anonymous):

But the answer sheet says 1/2. Where did I go wrong?

OpenStudy (hba):

You got \[\sqrt{2}/2\] =\[\sqrt{2}/\sqrt{2} \times \sqrt{2}\] So according to you the answer should be \[1/\sqrt{2}\]

OpenStudy (anonymous):

Correct.

OpenStudy (hba):

But now lets get to your question, \[f(x) = √(2x)\] f'(x)= Now think and derivate it correctly.

OpenStudy (anonymous):

f(x) =2(x)^1/2

OpenStudy (anonymous):

f'(x) = x^-1/2

OpenStudy (anonymous):

\[1/\sqrt{x}\]

OpenStudy (anonymous):

=\[\sqrt{x}/x\]

OpenStudy (anonymous):

So why is it not \[\sqrt{2}/2\] ?

OpenStudy (hba):

@Loushe You are not diffrentiating it right

OpenStudy (anonymous):

Chain rule?

OpenStudy (anonymous):

required here f'(x)=√2(1/2 *1/√x)=1/√(2x) hence f'(2)=1/√(2*2) =1/2

OpenStudy (hba):

@matricked Giving direct answers is not allowed Please check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

f(x) =(√2) *√x =(√2) *(x)^1/2

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

That helps me understand a lot.

OpenStudy (anonymous):

@hba but why to confse the person so much yes clear explanation is required

OpenStudy (hba):

@matricked As a rule of thumb, we would prefer the askers to participate in the solution process instead of having the answerers doing all the work. thats not really our concern; we are aimed at being a study site, not a free answering service. We would rather promote learning the material as opposed to just handing over answers. This social contract is aimed at curbing people from using the site to cheat with ...

OpenStudy (hba):

@matricked I hope you understand ?

OpenStudy (anonymous):

@hba i am not here to argue with the issue it is also upto the choice of askers how he learns and upto us how we make his learning easy ... if u r correct in ur ways then why so much botheration...

OpenStudy (hba):

@matricked These are the rules you have to follow them.

OpenStudy (anonymous):

@hba these medals aren't so improtant to me , neither i have asked anyone to click for the best response thanking u hope u understand ..

OpenStudy (hba):

@matricked This is not my concern whether you get 1000 medals on a question. The concern is that you must let the asker participate and not give direct answers.

OpenStudy (hba):

@matricked I hope this is clear to you ?

OpenStudy (anonymous):

@Loushe if you are clear with this question, then you can close this question...